What is the pOH of a solution containing an [H+] concentration of 2.75*10-3


1
Expert's answer
2021-05-05T01:41:58-0400

"pH=-log[H^+]=-log(2.75\times10^{-3})=2.561"

"pOH=14-pH=14-2.561=11.439"


Answer: 11.439